A professional photographer who specializes in wedding-related activities paid $45,000 for equipment that will have a $2000 salvage value after 5 years. He estimates that his costs associated with each event amount to $65 per day. If he charges $250 per day for his services, how many days per year must he be employed in order to break even at an interest rate of 10% per year?

The number of days required for break even is determined to be --------------- per year.
Answer #1

Solution:

We shall solve this using present worth analysis (other way Is using annual worth analysis)

Annual net revenue = revenue - cost

ANR = 250x - 65x = 185x ; where x is the number of days

We are given the initial cost as IC = $45,000; salvage value SV = $2,000 and number of years as N = 5

Interest rate is given as i = 0.10

Present worth, PW = -IC + ANR(P/A,i,N) + SV(P/F,i,N)

(P/A,i,N) = ((1+i)N-1)/i(1+i)N; so, (P/A, 0.10, 5) = ((1+0.10)5-1)/0.10*(1+0.10)5 = 0.61051/0.161051 = 3.79078

(P/F,i,N) = (1+i)-N so, (1 + 0.10)-5 = 0.62092

Breakeven occurs where present worth = 0, so

0 = - 45000 + 185x*3.79078 + 2000*0.62092

45000 - 1241.84 = 701.2943x

x = 43758.16/701.2943 = 62.396

Thus, the number of days required for break even is approximately 62 days.